# Rotational mass of anisotropic neutron stars within Rastall gravity

### v2

@@ 2.3 Anisotropy
Here $\zeta$ is a constant that represents the anisotropic stength, $m(r)$ is mass of the star at the distance $r$ measured from the center of the star, and $p$ is pressure at the distance $r$ from the center of the star. It is obvious that near the surface of the star, the anisotropy is proportional to the compactness $M/R$. The model showed by Eq. (40) is an ansatz which is assumed to be proportional to the radial pressure and to the compactness. A paramount feature of the ansatz is that the compactness goes as $r^{2}$ when $r\rightarrow 0$ securing the required vanishing of the pressure anisotropy at the center of symmetry. Another motivation for this ansatz was drawn from astrophysical consideration that in the non-relativistic regime where the compactness is much smaller than unity, the anisotropy of pressure can be expected not to play an important role. Furthermore, when the tangential pressure exceeds the radial pressure (i.e. $\zeta<0$), the anisotropy may be more efficient in supporting highly compact bodies against gravitational collapse than in the opposite case (i.e. $\zeta>0$) [67].
@@ 3 Numerical results and discussion
Another important differential equation is Eq. (21), which contributes in giving the value of factor $\bar{\omega}$ and $\frac{d\bar{\omega}}{dr}$ in other differential equations. Since the value of $\bar{\omega}_{c}$ as the initial value of $\bar{\omega}$ can be chosen arbitrarily, we choose 240 $s^{-1}$ in this work. It has to be highlighted that this value is much less than the value used in Ref. [4], i.e. $\bar{\omega}$ = 3000 s-1. Choosing our value of $\bar{\omega}_{c}$ is also relatively wise since it is more suitable for HT formalism. We have to consider that $\bar{\omega}(r)$ increases as $r$ increases. Although we have a fact that according to Ref. [14], the HT formalism is accurate to better than 1 per cent even for the fastest millisecond pulsars, putting $\bar{\omega}_{c}$ = 3000 s-1 in the calculation is potential to be awful, for the HT formalism is basically constructed for slowly rotating relativistic stars. So our value of $\bar{\omega}_{c}$ is safer than the one in Ref. [4]. Comparing to Ref. [7] and Ref. [16] which chose $\bar{\omega}_{c}=0$, our choice of $\bar{\omega}_{c}$ is also more realistic for general rotating NS, for their choice tends to only agree the NSs with very slow rotation.
Fig. 1 shows the relationship between radius and rotational mass of anisotropic neutron stars (NSs). The figure clearly demonstrates that Rastall gravity differs from GR ($\lambda=0$). The Rastall parameter $\lambda$, can either increase or decrease the mass of the NS within specific ranges of $R$. In this study, we use three values of $\lambda$: $\lambda=0.00019$, $\lambda=0.00038$, and $\lambda=0.00071$.
@@ Acknowledgements
MLP gratefully acknowledges the Indonesia Endowment Fund for Education (LPDP) for financial support. Special thanks are extended to Anna Campoy Ordaz for making her code publicly available, which allowed MLP to modify it for use in this work. FPZ would like to thank to the Ministry of Education, Culture, Research, and Technology of the Republic of Indonesia (Kemendikbudristek) through the LPPM ITB for partial financial support. HLP acknowledges the Ganesha Talent Assistantship Institut Teknologi Bandung for financial support. MLP and HLP would also like to thank the members of the Theoretical Physics Groups at Institut Teknologi Bandung for their hospitality. MFARS is supported by the Second Century Fund (C2F), Chulalongkorn University, Thailand.

### v3

@@ 2.4 Anisotropy
Here $\zeta$ is a constant that represents the anisotropic stength. When the tangential pressure exceeds the radial pressure (i.e. $\zeta<0$), the anisotropy may be more efficient in supporting highly compact bodies against gravitational collapse than in the opposite case (i.e. $\zeta>0$) [67].
@@ 3 Numerical results and discussion
Another important differential equation is Eq. (LABEL:omega), which contributes in giving the value of factor $\bar{\omega}$ and $\frac{d\bar{\omega}}{dr}$ in other differential equations. Since the value of $\bar{\omega}_{c}$ as the initial value of $\bar{\omega}$ can be chosen arbitrarily, we choose 240 $s^{-1}$ in this work. It has to be highlighted that this value is much less than the value used in Ref. [4], i.e. $\bar{\omega}$ = 3000 s-1. Choosing our value of $\bar{\omega}_{c}$ is also relatively wise since it is more suitable for HT formalism. We have to consider that $\bar{\omega}(r)$ increases as $r$ increases. Although we have a fact that according to Ref. [14], the HT formalism is accurate to better than 1 per cent even for the fastest millisecond pulsars, putting $\bar{\omega}_{c}$ = 3000 s-1 in the calculation is potential to be awful, for the HT formalism is basically constructed for slowly rotating relativistic stars. So our value of $\bar{\omega}_{c}$ is safer than the one in Ref. [4]. Comparing to Ref. [7] and Ref. [16] which chose $\bar{\omega}_{c}=0$, our choice of $\bar{\omega}_{c}$ is also more realistic for general rotating NS, for their choice tends to only agree the NSs with very slow rotation.
Fig. 1 shows the relationship between radius and rotational mass of anisotropic neutron stars (NSs). The figure clearly demonstrates that Rastall gravity differs from GR ($\lambda=0$). The Rastall parameter $\lambda$, can either increase or decrease the mass of the NS within specific ranges of $R$. In this study, we use three values of $\lambda$: $\lambda=0.00019$, $\lambda=0.00038$, and $\lambda=0.00071$. Table 3 displays the calculated values of $\kappa\lambda$ and $\Lambda$ for each value of $\lambda$. It is verified that the values of $\lambda$ used in this work do not violate any restrictions specified in Eqs. (4), (7), (8), (9), and (10). Thus, all values of $\lambda$ employed in this work are allowed.
$\lambda$ $\kappa\lambda$ $\Lambda$
0.00019 0.08493 0.12862
0.00038 0.08652 0.13230
0.00071 0.08928 0.13888
Table 3: The values of $\lambda$ and the corresponding values of $\kappa\lambda$ and $\Lambda$.
@@ Acknowledgements
We thank Anna Campoy Ordaz for making her code publicly available, which allowed MLP to modify it for use in this work. MLP gratefully acknowledges the Indonesia Endowment Fund for Education (LPDP) for financial support. FPZ would like to thank to the Ministry of Higher Education, Research, and Technology of the Republic of Indonesia (Kemendiktisaintek) through the LPPM ITB for partial financial support. MLP and HLP would also like to thank the members of the Theoretical Physics Groups at Institut Teknologi Bandung for their hospitality. MFARS is supported by the Second Century Fund (C2F), Chulalongkorn University, Thailand.